OpenStudy (anonymous):

square root of 10 open parenthesis 4 minus square root of 5 close parenthesis plus square root of 8 open parentheses square root of 5 end root plus 2

OpenStudy (radar):

\[\sqrt{10}(4-\sqrt{5})+\sqrt{8}(\sqrt{5}+2)\]Is this the problem?

OpenStudy (anonymous):

yes! :)

OpenStudy (radar):

\[4\sqrt{10}-\sqrt{50}+\sqrt{40}+2\sqrt{8}\]\[4\sqrt{10}-5\sqrt{2}+2\sqrt{10}+4\sqrt{2}\]\[6\sqrt{10}-\sqrt{2}\]

OpenStudy (anonymous):

thank you ! makes sense now